OpenStudy (anonymous):

Solve the given differential equation using the proper substitution: dy/dx=(x+y+1)^2 Someone...anyone...?

OpenStudy (anonymous):

can i ask what is proper substitution ?

OpenStudy (turingtest):

wolfram makes good use of \[v(x)=x+y(x)\]

OpenStudy (anonymous):

u can do it with \(z=x+y+1\) but i dont know...is this proper substitution?

OpenStudy (anonymous):

it starts out as u=x+y+1 and then derives u to get du=1+dy/dx. After this it becomes dx=(du/1+u^2) and I can follow the algebra; I'm just confused as to what's going on and why.

OpenStudy (amistre64):

your changeing the function into hopefully an easier problem thru the substitution

OpenStudy (amistre64):

change of variables is another phrase ive seen used

OpenStudy (anonymous):

Hmm...I need to see this kind of problem solved so that I can see the process I think. I've had trouble finding one similar online. I have the solutions manual but I need to understand what's going on and there are several steps which are skipped.

OpenStudy (anonymous):

Perhaps I should continue my search...

OpenStudy (amistre64):

the first setup: dy/dx tells us that y is a function of x we want to find an easier setup for us to work with and solve for u = x + y(x) + 1 ; derive it to determine a suitable du/dx, and possibly a dy/du since:\[\frac{dy}{dx}=\frac{dy}{du}\frac{du}{dx}\] du/dx = dx/dx + dy/dx + 0 = 1 + dy/dx yeah, it prolly help to keep searching :)

OpenStudy (anonymous):

Yeah, I'm on the verge of understanding here. I'm not quite sure why the solution goes on to simplify in terms of dx though. It reads dx=du/(1+u^2) and then integrates this. I'm confused as to why and what is happening here. I shall continue searching. Thank you kindly for your help though! :)

OpenStudy (amistre64):

\[\frac{du}{dx}=1+\frac{dy}{dx}\to\ \frac{dy}{dx}=\frac{du}{dx}-1\] using this then \[\frac{dy}{dx}=u^2\] \[\frac{du}{dx}-1=u^2\] \[\frac{du}{dx}=u^2+1\]

OpenStudy (anonymous):

Yes, this is exactly what the solution is giving me. Why are we solving for dx by itself though?

OpenStudy (amistre64):

by separating this, we can solve for du \[\frac{du}{u^2+1}=dx\] \[\int \left(\frac{du}{u^2+1}=dx\right)\] \[\int \frac{du}{u^2+1}=\int dx\]

OpenStudy (anonymous):

Okay I see...

OpenStudy (amistre64):

now, since we know that : u = x+y+1 once we find a suitable "u" in terms of x, we can solve for y y = u-x-1

OpenStudy (anonymous):

wait wait wait...why all of a sudden are they u-x-1 rather than u+x+1? Why did the signs change?

OpenStudy (amistre64):

remember how we defined the substitution: u = x + y + 1 ; but we dont want "u", we want "y" what is y

OpenStudy (amistre64):

or rather:\[u(x)=x+y(x)+1\]define \(y(x)\) \[y(x)=u(x)-x-1\]

OpenStudy (anonymous):

Ahhhh...so jump back a step though? Are we not integrating both sides of the equation where we separated du/(1+u^2) and dx?

OpenStudy (amistre64):

we are

OpenStudy (amistre64):

we are determining as solution to u(x)

OpenStudy (anonymous):

ohhh so then I get \[\tan^{-1} (x+c)-x-1\]

OpenStudy (anonymous):

I think I get it.

OpenStudy (amistre64):

it looks good to me; its easier prolly to think in terms of functions in stead of variables to get your head around this

OpenStudy (anonymous):

Yes I think you've got a point there. I just finished solving Bernoulli DEs and thought these looked easier but for some reason the substitution part threw me a curve ball. What I find kind of confusing is that the original DE had everything in the brackets squared. \[\left( x+y+1 \right)^{2}\] Where did the power or two disappear to? Was it eaten up in the substitution as well?

OpenStudy (amistre64):

it did not disappear, we used it in the process (x+y(x)+1)^2 ; such that u(x) = x+y(x)+1, therefore (x+y(x)+1)^2 = \(u^2(x)\)

OpenStudy (anonymous):

Ohhh so THAT's why we had

OpenStudy (anonymous):

u^2 in the earlier steps. It's all making sense now. :D

OpenStudy (anonymous):

Thank you very very much! You've been a great help!

OpenStudy (amistre64):

youre welcome, and good luck ;)

OpenStudy (anonymous):

Thanks, cheers!